OpenStudy (anonymous):

a ball rolls off a horizontal shelf a height h above the floor and takes 0.5s to hit.For the ball to take 1.0s to reach the floor the shelf's height above the floor would have to be A) 2h B)3h C) sqrt(2) h D)4h

OpenStudy (anonymous):

rolls off=no initial vertical component of velocity. \(S=0t+\frac{1}{2}(9.8) t^2\) \(H_1=4.9t_1^2\) when \(t_2=2t_1\)(i.e. 1s is double 0.5s) \(H_2=4.9t_2^2=4.9(2t_1)^2=4.9(4t_1^2)=4(4.9t_1^2)=4H_1\)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you're welcome :)